# Representation Learning in a Decomposed Encoder Design for Bio-inspired Hebbian Learning

### v1

@@ 4.2 Experimental Results
4.3 t-SNE Visualization
We qualitatively support our empirical classification results (Table 1) by plotting the t-SNE Van der Maaten and Hinton (2008) embedding of the encodings on CLAPP and on our method with decomposition (Figure 2). The plot shows the potential of our decomposition approach in separating representations of different classes. However, we note that separability of t-SNE clusters depends on the chosen dataset and the visualized classes. For instance, the class with speed limit sign 20km/h will form a cluster that is much closer to classes of other speed limit signs. Further visualization on different classes as well as on STL10 dataset can be found in the appendix.
4.4 Ablation Study
@@ 4.5 Multi-Target Classification
To further understand how of the decomposition approach scales to real world applications, we benchmark our method on the CODEBRIM dataset Mundt et al. (2019). We train our encoders in a similar manner to previous experiments. After that, we freeze the weights of the encoder and train a linear downstream multi-target classifier on representations created by the frozen encoder (Table 3).
@@
MetaQNN-1 Mundt et al. (2019) $66.2\pm 1.6$ $87.6\pm 0.5$
HingeCPC Illing et al. (2021); Oord et al. (2018) $33.87\pm 0.12$ $82.27\pm 0.27$
CLAPP Illing et al. (2021) $37.78\pm 0.5$ $83.05\pm 0.6$
@@
Moreover, we observe that the pipeline with the local learning slightly outperforms the pipeline trained with backpropagation. One possible explanation is that the CODEBRIM dataset is a much smaller dataset than STL10 or GTSRB. As observed in previous works Lagani et al. (2021), training with Hebbian rules can in some cases be more sample efficient.
4.6 CODEBRIM with Perturbations
Figure 3: Box plot illustration to summarize the multi-target test classification results under different perturbation settings. Multi-target classification accuracy (%) refers to classification of all classes correctly at the same time. Each box plot shows the median, first and third quartile and minimum and maximum values obtained from all experimental results under different perturbation settings. Further results for each type of perturbation can be found in the appendix.
In this section, we evaluate the robustness of the proposed models to common perturbations. We choose perturbations that can be encountered in real applications. Gaussian noise may appear due to poor illumination or overheated sensors. Motion blur is caused by motion of the camera during acquisition. Planckian Jitter creates realistic physics-based variations in chromaticity Zini et al. (2022). Shadows and brightness artifacts can occur when capturing data in the wild in the presence of some light source.
We simulate these perturbations on the CODEBRIM data with the Kornia package Riba et al. (2020) and study how the performance degrades across different models, presented in the previous section. Note that our encoders and their corresponding classifiers were not re-trained with these additional perturbations as data augmentations. The models are exposed to these perturbations for the first time during inference on the test set. Our goal is to show how the addition of multiple visual invariant operators can help to mitigate the performance degradation on adversarial examples.
Figure 3 summarizes the multi-target downstream classification performance for different perturbations with different intensities. It appears that overall, decomposition leads to better classification performance on adversarial examples both in the case of the biologically plausible model and the backpropagation model.
4.7 Experimental Results on Video Streams
Figure 4: Diagram illustrating the training procedure for motion sensitive neural network (MovNet) on video data. See appendix for further details.
For video data, we extend our framework with another encoder (MovNet) for encoding motion representations. Inspired by previous works on motion representation learning(Mineault et al. (2021); Rideaux and Welchman (2020); Huang et al. (2021)), we propose a self-supervised task to learn motion features from a sequence of frames. Figure 4 illustrates the training procedure. The task is based on the discrimination between egomotion and motion of objects in the scene on one hand, and the prediction of the speed of motion on the other. We hypothesize that if an agent is capable of differentiating between the camera motion and the motion of objects, then it should learn to extract latent representations encoding motion information, which can then be used for a better and more robust action and object recognition in video streams. Egomotion is generated by sampling a single video frame randomly and generating a sequence of overlapping patches. This encoder is trained end-to-end with backpropagation and is included in our experiments to further illustrate how a decomposed design for representation learning can be beneficial in achieving better performance. For further implementation details, see appendix.
Method Top-1 Acc Top-5 Acc
Supervised VGG $45.3\pm 0.5$ $77.2\pm 0.3$
HingeCPC Illing et al. (2021); Oord et al. (2018) $33.7\pm 0.18$ $60.43\pm 0.8$
CLAPP Illing et al. (2021) $29.66\pm 0.3$ $53.81\pm 0.7$
[…]
$RGNorm$+HingeCPC $3.74\pm 1.1$ $14.83\pm 1.4$
$DTCWT$+HingeCPC $26.32\pm 0.86$ $49.19\pm 0.42$
[…]
$RGNorm$+CLAPP $5.66\pm 1.33$ $17.5\pm 1.1$
$DTCWT$+CLAPP $28.7\pm 0.4$ $50.8\pm 0.8$
[…]
ours(Local) $35.41\pm 0.7$ $62.8\pm 0.6$
[…]
ours(Local)+MovNet $38.88\pm 0.3$ $66.28\pm 0.6$
Table 4: Performance comparison in terms of classification accuracy (%) on UCF101 test set. The best performing self-supervised model is highlighted in red, in bold the best performing model for each category: baselines, models trained with HingeCPC, models trained with CLAPP, decomposed encoders, decomposed encoders with motion pipeline.
The results for the action recognition downstream task are presented in Table 4. The addition of a single decomposition operator leads to worse performance compared to the network without any operators. This contrasts with the self-supervised results for object recognition in previous sections where the inductive bias provided by at least one decomposition operator helps the network learn more linearly separable representation (LBP on GTSRB and DTCWT on STL10 and COBEBRIM). But despite the fact that each single encoder of our framework has a worse performance than the encoder without any visual operator, the overall framework achieves the best classification accuracy. The proposed combination of operators is flexible enough to achieve better performance compared to baseline encoders.
The framework trained with local learning rules shows more substantial improvement compared to its baseline ($6\%$ improvement of ours(Local) compared to CLAPP and $2\%$ improvement for ours(Backprop) compared to HingeCPC). This is consistent with results from the previous sections. Furthermore, adding the encoder trained on motion information to our framework leads to even better performance.
The addition of a motion sensitive module leads to a more linearly separable latent representation as suggested by the better classification performance. These results indicate that motion information can be helpful in action recognition from video data. This also supports our hypothesis that decomposition into multiple parallel representations can alleviate representation learning in different application contexts and on different data forms.
@@ 5 Conclusion
In this manuscript, we presented a framework to study the impact of decomposition as an inductive bias on representation learning with local plasticity rules as well as backpropagation. Our experiments indicate that the decomposition of an input signal with well understood image transformation operators can be a powerful tool to learn more generalizable latent representations. Additionally, it can be beneficial to close the gap between representation learning with local rules and backpropagation.
These results showcase the importance of domain knowledge and application context understanding when choosing the decomposition pipeline as each dataset requires different quasi-invariances. Formalization of the choice of operators is an important issue to have stable performance across different domains. A single decomposition operator projects the input signal into a quasi-invariant space and gets rid of some variations of the input which might prove useful to solve a downstream task. Hence, it is crucial that the chosen operators are complementary so that the whole system is able to achieve better performance in different contexts.

### v2

@@ 4.2 Experimental Results
We qualitatively support our empirical classification results (Table 1) by plotting the t-SNE [30] embedding of the encodings on CLAPP and on our method with decomposition (Figure 2). The plot shows the potential of our decomposition approach in separating representations of different classes. However, we note that separability of t-SNE clusters depends on the chosen dataset and the visualized classes. For instance, the class with speed limit sign 20km/h will form a cluster that is much closer to classes of other speed limit signs.
4.3 Ablations
@@ 4.4 Multi-Target Classification
To further understand how of the decomposition approach scales to real world applications, we benchmark our method on the CODEBRIM dataset [35]. We train our encoders in a similar manner to previous experiments. After that, we freeze the weights of the encoder and train a linear downstream multi-target classifier on representations created by the frozen encoder (Table 3).
@@
MetaQNN-1 [35] $66.2\pm 1.6$ $87.6\pm 0.5$
HingeCPC [18, 39] $33.87\pm 0.12$ $82.27\pm 0.27$
CLAPP [18] $37.78\pm 0.5$ $83.05\pm 0.6$
@@
Moreover, we observe that the pipeline with the local learning slightly outperforms the pipeline trained with backpropagation. One possible explanation is that the CODEBRIM dataset is a much smaller dataset than STL10 or GTSRB. As observed in previous works [23], training with Hebbian rules can in some cases be more sample efficient. Further results in the appendix show that our decomposed achieves better performance on adversarial and noisy examples.
@@ 5 Conclusion
In this manuscript, we presented a framework to study the impact of decomposition on representation learning with local plasticity rules and backpropagation. Our experiments indicate that decomposing input signals with well-understood image transformation operators can enhance the generalizability of latent representations and narrow the gap between local learning rules and backpropagation.
These results underscore the importance of domain knowledge and application context in choosing the decomposition pipeline, as different datasets require different quasi-invariances. Formalizing the choice of operators is a crucial next step for stable performance across domains. Complementary operators are essential to project input signals into complementary quasi-invariant spaces, improving performance in various contexts.